OpenStudy (anonymous):

1/1+ sin 0 + 1/ 1- sin 0 A. 2 cos2 0 B. 2 sec2 0 C. 2 csc2 0 D. 2 cot2 0

OpenStudy (loser66):

\[\frac{1}{1+sinx}+\frac{1}{1-sinx}?\]

OpenStudy (loser66):

\[i~replace~\theta=x\]

OpenStudy (anonymous):

Ok im really sorry but i have no idea at all about algebra :(

OpenStudy (anonymous):

ok i get what you did now but what would be my next step?

OpenStudy (loser66):

make them have the same denominator

OpenStudy (anonymous):

So wouldnt it be A

OpenStudy (loser66):

no, you have to do step by step. i didn't get it yet. but if you do, I can check.

OpenStudy (anonymous):

I dont even know what steps 2 take. thats my biggest problem with algebra i just dont get it at all..

OpenStudy (anonymous):

They already have the same Denominator besides that one is plus and the other is minus?

OpenStudy (loser66):